Problem 2 Your company, Dominant Retailer, Inc., is considering a project whose data are shown below. Revenue and cash operating expenses are expected to be constant over the project's 5 year expected operating life; annual sales revenue is $95,000.00 and cash operating expenses are $37,500.00. The new equipment's cost and depreciable basis is $135,000.00 and it will be depreciated by MACRS as 5 year property. The new equipment replaces older equipment that is fully depreciated but can be sold for $7,500. In addition, the new equipment requires an additional $5,000 of net operating working capital, which can be fully recovered at the end of the project. The new equipment is expected to be sold for $10,995 at the end of the project in year 5. The marginal tax rate is 28.00%. What is the project's Initial Cash Outlay at Year 0? Note: Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.

Using the information from problem 2 on Dominant Retailer, Inc., what is the NPV of the Project if Dominant Retailer’s WACC is 12.75%? Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.
Answer #1

Calculating projects initial outlay of project in year 0

Cost of new equipment = $135000, Initial investment in working capital = $5000

Salvage value of old equipment = 7500, Book value of old equipment = 0

Net proceeds from sale of old equipment = Salvage value of old equipment - Tax on gain from sale of old equipment = Salvage value of old equipment - tax rate(Salvage value of old equipment - Book value of old equipment) = 7500 - 28%(7500 - 0) = 7500 - 2100 = 5400

Initial outlay in year 0 = Cost of new equipment + Initial investment in working capital - Net proceeds from sale of old equipment = 135000 + 5000 - 5400 = 134600

Projects initial outlay in year 0 = 134600.00

Calculating NPV of the project

Depreciation for a year under 5 year MACRS = Rate of depreciation for a year x Cost of equipment

Rate of depreciation for year 1 under 5 year MACRS = 20%

Depreciation for year 1 under 5 year MACRS = Rate of depreciation for year 1 under 5 year MACRS x 135000 = 27000

Similarly depreciation can be calculated for other years. Depreciation schedule is given in the table below

Depreciation under 5 year MACRS class
Year 1 2 3 4 5
Depreciation rate 20.00% 32.00% 19.20% 11.52% 11.52%
Depreciation 27000.00 43200.00 25920.00 15552.00 15552.00

After tax operating cash flow for a year = EBIT(1- tax rate) + Depreciation = (Revenue - expenses - depreciation)(1-tax rate) + depreciation

After tax operating cash flow for year 1 = (95000 - 37500 - 27000)(1-28%) + 27000 = 30500 x 72% + 27000 = 21960 + 27000 = 48960

Similarly we can calculate the after tax operating cash flow for year 2 to year 5, we get the following table of cash flows

Calculating After tax Operating Cash flow
Year 1 2 3 4 5
Revenue 95000.0000 95000.0000 95000.0000 95000.0000 95000.0000
Operating costs 37500.0000 37500.0000 37500.0000 37500.0000 37500.0000
Depreciation 27000.0000 43200.0000 25920.0000 15552.0000 15552.0000
EBIT 30500.0000 14300.0000 31580.0000 41948.0000 41948.0000
EBIT(1-tax rate) 21960.0000 10296.0000 22737.6000 30202.5600 30202.5600
Plus: Depreciation 27000.0000 43200.0000 25920.0000 15552.0000 15552.0000
After tax operating Cash Flow 48960.0000 53496.0000 48657.6000 45754.5600 45754.5600

Book value of new equipment at end of year 5 = Cost of new equipment - sum of depreciation for year 1 to 5 = 135000 - (27000 + 43200 + 25920 + 15552 + 15552) = 135000 - 127224 = 7776

Terminal cash flow in year 5 = Salvage value of new equipment at end of 5 years - tax from gain on sale of new equipment + Recovery of investment in net working capital

= Salvage value of new equipment at end of 5 years - tax rate(Salvage value of new equipment at end of 5 years - Book value of new equipment at end of 5 years)- + Recovery of investment in net working capital

= 10995 - 28%(10995 - 7776) + 5000 = 10995 - 901.32 + 5000 = 15093.68

Net present value of project = - Initial outlay in year 0 + Sum of present values of after tax cash flows discounted at WACC + Present value of terminal cash flow discounted at WACC

= - 134600 + 48960 / (1 + 12.75%) + 53496 / (1 + 12.75%)2 + 48657.60 / (1 + 12.75%)3 + 45754.56 / (1 + 12.75%)4 + 45754.56 / (1 + 12.75%)5 + 15093.68 / (1 + 12.75%)5

= -134600 + 43423.5033 + 42081.2090 + 33946.9712 + 28311.8455 + 25110.2843 + 8283.4715 = 46557.2839 = 46557.28 (rounded to two places off decimal)

Hence NPV of the project = 43557.28